OpenStudy (anonymous):

Find the Taylor series of f(x)=ln(1+2x) around x_0=1

OpenStudy (anonymous):

this doesn't look fun around \(1\) ?

OpenStudy (anonymous):

constant is \(f(1)=\ln(3)\)

OpenStudy (anonymous):

the derivative is \[\frac{2}{1+2x}\] at \(1\) you get \(\frac{2}{3}\) so the next term is \[\frac{2}{3}(x-1)\]

OpenStudy (anonymous):

how do we write the taylor series